OpenStudy (anonymous):

Evaluate the series 9 + 3 + 1 +... pleaseee help me!! i dont know how to do this:)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

it is a geometric progression with common ratio 1/3 Sum = a/(1-r) = 9/(1-1/3) = 27/2

OpenStudy (anonymous):

like so its a multiple choise qustion and they are.. a. 13 b.3 c.27 d.1.5

OpenStudy (anonymous):

13.5 is the answer

OpenStudy (anonymous):

ahhhh thanks

OpenStudy (anonymous):

no it is 27 duh1

OpenStudy (amistre64):

9 + 3 + 1 +... an infinite series eh ...

OpenStudy (amistre64):

well; the /3 is a good assumption given the limited data ....